If it would take one printer 4 hours to print the company newsletter and it would take a different printer 6 hours to do the same job, how long would it take if both printers are used?
Answer by stanbon(75887)

If it would take one printer 4 hours to print the company newsletter and it would take a different printer 6 hours to do the same job, how long would it take if both printers are used?
======
1st printer DATA:
time = 4 hrs/job ; rate = 1/4 job/hr
---
2nd printer DATA:
time = 6 hrs/job ; rate = 1/6 job/hr
----
Together DATA:
time = x hrs/job ; rate = 1/x job/hr
-----------------------------
Equation:
rate + rate = together rate
1/4 + 1/6 = 1/x
-------
6x + 4x = 4*6
----
10x = 24
x = 2.4 hrs = 2 hrs 24 min (time to do the job together)
--------------
Cheers,
Stan H.
--------------